# Regression-Based Proximal Causal Inference

### v1

@@ A.3.3 Extension to Polytomous Regression
$=\tilde{\beta}_{0t}+\beta_{at}^{*}A+\beta_{ut}^{*}\sum_{k=1}^{K}\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y=1))+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\ \textrm{, where $\beta_{at}^{*}=\beta_{at}$.}$
$S=\sum_{k=1}^{K}\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y=1))=\sum_{k=1}^{K}\left(\tilde{\alpha}_{0k}+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}+\alpha_{uk}m(A,Z)\right)\propto_{L}m(A,Z)$
@@
    Compute $\hat{S}=\sum_{k=1}^{K}\mathrm{logit}(\widehat{\mathrm{Pr}}(W=k|A,Z,Y=1))=\sum_{k=1}^{K}\left(\hat{\alpha}_{0k}^{*}+\hat{\alpha}_{ak}^{*}A+\hat{\alpha}_{zk}^{*}Z+\sum_{t=1}^{T}\hat{\alpha}_{ykt}^{*}\right)$ using the estimated coefficients from the first-stage regression.
@@
Similarly, following steps in Section A.3.2, we can have interaction terms involved in regression function specification. Details are omitted.
A.4 Identity and Log Links Combination
In this section, we present details on the continuous $Y$ and count $W$ and count $Y$ and continuous $Y$ cases, which correspond to the gray cells in the table below.
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link)
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link) Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link)
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
A.4.1 Log $Y$ Link and Identity $W$ Link
@@
$U=E[U|A,Z]+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z);$ (76)
$*.\displaystyle|E[U|A,Z]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of$\epsilon$ exists for any real value *.}$.
@@
Due to the shared collapsibility of identity and log links, iterating similar steps in Sections A.1 and A.2 leads to the answer. We only provide minimal algebraic details.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.,
@@
A.4.2 Identity $Y$ Link and Log $W$ Link
@@
$U=E[U|A,Z]+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z);$
$*.\displaystyle|E[U|A,Z]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of$\epsilon$ exists for any real value *.}$.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.,
@@
A.5 Identity and Logit Links Combination
In this section, we present details on the continuous $Y$ and binary $W$ and binary $Y$ and continuous $Y$ cases, which correspond to the gray cells in the table below.
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link)
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link)
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link) Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
@@ A.5 Identity and Logit Links Combination
A.5.1 Logit $Y$ Link and Identity $W$ Link Combination with No Interaction Term
@@
$U=\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(Y=0);$ (79)
$*.\displaystyle|E[U|A,Z,Y=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.}$.
@@
where the second equality holds from (83). We can find that $E[W|A,Z,Y=0]$ and $E[W|A,Z,Y=1]$ have the same form except for the intercept term. Therefore, we can combine (81) and (84), and write them into a single representation:
$E[W|A,Z,Y]=\alpha_{0}+\alpha_{u}E[U|A,Z,Y]+\alpha_{y}Y=\alpha_{0}+\alpha_{u}t(A,Z)+\tilde{\alpha}_{y}Y$
where $\textrm{where }\tilde{\alpha}_{y}=\alpha_{y}+\alpha_{u}c$
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=0]=t(A,Z)$, i.e.,
@@
A.5.2 Logit $Y$ Link and Identity $W$ Link Combination with Interaction Terms
@@
$U=\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(Y=0);$ (87)
$*.\displaystyle|E[U|A,Z,Y=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.}$.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=1]=t(A,Z)$, i.e.,
@@
A.5.3 Identity $Y$ Link and Logit $W$ Link Combination with No Interaction Term
@@
$U=\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(W=0);$
$*.\displaystyle|E[U|A,Z,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.}$.
@@
A.5.4 Identity $Y$ Link and Logit $W$ Link Combination with Interaction Terms
@@
$U=\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(W=0);$
$*.\displaystyle|E[U|A,Z,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.}$.
@@
$E[Y|A,Z,W]=\beta_{0}+\beta_{a}A+\beta_{u}h(A,Z)+\beta_{w}W+\beta_{uw}h(A,Z)W$
@@
A.6 Log and Logit Links Combination
In this section, we present details on the count $Y$ and binary $W$ and binary $Y$ and count $Y$ cases, which correspond to the gray cells in the table below.
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link)
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link)
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link)
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
It is no surprise that everything in Section A.5 extends here due to similar properties of the identity and the log links. The only differences are replacing the identity link from Section A.5 and the intercept terms absorbing constants from MGFs; a similar phenomenon has been discussed in Section A.2. Due to algebraic similarities, we omit the details and directly state the results.
A.6.1 Logit $Y$ Link and Log $W$ Link Combination with No Interaction Term
@@
$U=\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(Y=0);$
$*.\displaystyle|E[U|A,Z,Y=0]|<\infty,\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.}$.
@@
Following algebra in Section A.2 and A.5.1, we obtain the following two regression equations which imply Result 12:
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=0]=t(A,Z)$, i.e.,
@@
A.6.2 Logit $Y$ Link and Log $W$ Link Combination with Interaction Terms
@@
$U=\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(Y=0);$
$*.\displaystyle|E[U|A,Z,Y=0]|<\infty,\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.}$.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=0]=t(A,Z)$, i.e.,
@@
A.6.3 Log $Y$ Link and Logit $W$ Link Combination with No Interaction Term

### v3

@@ A.4.3 Extension to Polytomous Regression
$=\beta_{0t}^{*}+\beta_{at}^{*}A+\sum_{k=1}^{K}\beta_{utk}^{*}\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y=0))+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\ \textrm{, where $\beta_{at}^{*}=\beta_{at}$.}$
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,W=0,Y=0]=m(A,Z)$ where each $S_{k}$ is a linear combination of the element $m_{v}(A,Z)$ of $m(A,Z)$, i.e.
$S=(S_{1},\ldots,S_{K})$
$m(A,Z)=(m_{1}(A,Z),\ldots,m_{V}(A,Z))$
$S_{k}=\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y=0))=\tilde{\alpha}_{0k}+\alpha_{uk}^{*^{\prime}}m(A,Z)\propto_{L}m(A,Z)$
$S=Cm(A,Z),\ where\ C\in R^{K\times V}\ of\ rank\ V$
Note that when $(W,Z)$ has a higher dimension than $U$, i.e., $\min(T,K)>V$, the rows of $C$ will not be linearly independent, leading to collinearity among the elements of $S$. This collinearity will be reflected in the second-stage regression. Therefore, if the estimated variance of the $S$ coefficients in the second-stage regression is excessively large, one possible explanation is that $U$ may be of significantly lower dimension than $W$ and $Z$.
@@
    Specify a linear regression function for $\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y))$. Say $W\sim A+Z+Y$ for all strata of $W$.
@@
    Compute $\hat{S}=\left(\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=0)),\ldots,\mathrm{logit}(\mathrm{Pr}(W=K|A,Z,Y=0))\right)$ using the estimated coefficients from the first-stage regression.
@@
Similarly, following steps in Section A.4.2, we can have interaction terms involved in regression function specification. Moreover, since the proof above does not require $K=T$, similar procedures apply to the cases where $(Y,W)$ are binary and multinomial (more than two categories), respectively, which correspond to the two gray cells of the table below. Details are omitted.
A.5 Identity and Log Links Combination
In this section, we present details on the continuous $Y$ and count $W$ and count $Y$ and continuous $Y$ case, which correspond to the gray cells in the table below.
A.5.1 Log $Y$ Link and Identity $W$ Link
@@
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon$ exists for any real value *.
@@
Due to the shared collapsibility of identity and log links, iterating similar steps in Sections A.2 and A.3 leads to the answer. We only provide minimal algebraic details.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.
@@
A.5.2 Identity $Y$ Link and Log $W$ Link
@@
$U|A,Z\sim E[U|A,Z]+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z);$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon$ exists for any real value *.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.
@@
A.6 Identity and Logit Links Combination
In this section, we present details on the continuous $Y$ and binary $W$ and binary $Y$ and continuous $Y$ case, which corresponds to the gray cells in the table below.
@@ A.6 Identity and Logit Links Combination
A.6.1 Logit $Y$ Link and Identity $W$ Link Combination with No Interaction Term
@@
$U|A,Z,Y=0\sim\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|Y=0;$ (81)
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.
@@
$\mathrm{logit}(\mathrm{Pr}(Y=1|A,Z))=\beta_{0}^{*}+\beta_{a}^{*}A+\beta_{u}^{*}S$
@@
where the second equality holds from (85). We can find that $E[W|A,Z,Y=0]$ and $E[W|A,Z,Y=1]$ have the same form except for the intercept term. Therefore, we can combine (83) and (86), and write
$E[W|A,Z,Y]=\alpha_{0}+\alpha_{u}E[U|A,Z,Y]+\alpha_{y}Y=\alpha_{0}+\alpha_{u}t(A,Z)+\tilde{\alpha}_{y}Y,$
where $\textrm{where }\tilde{\alpha}_{y}=\alpha_{y}+\alpha_{u}c.$
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=0]=t(A,Z)$, i.e.
@@
A.6.2 Logit $Y$ Link and Identity $W$ Link Combination with Interaction Terms
@@
$U|A,Z,Y=0\sim\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|Y=0;$ (89)
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=1]=t(A,Z)$, i.e.
@@
A.6.3 Identity $Y$ Link and Logit $W$ Link Combination with No Interaction Term
@@
$U|A,Z,W=0\sim\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|W=0;$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.
@@
A.6.4 Identity $Y$ Link and Logit $W$ Link Combination with Interaction Terms
@@
$U|A,Z,W=0\sim\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|W=0;$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.
@@
$E[Y|A,Z,W]=\beta_{0}+\beta_{a}A+\beta_{u}h(A,Z)+\tilde{\beta}_{w}W+\beta_{uw}h(A,Z)W$
@@
By similar algebra steps, we can solve the case where the logit link is applied to the multinomial $Y$ or $W$. The procedures are given in the two gray cells of the table below. Details are omitted.
A.7 Log and Logit Links Combination
In this section, we present details on the count $Y$ and binary $W$ and binary $Y$ and count $Y$ case, which corresponds to the gray cells in the table below.
It is not entirely surprising that the steps followed in Section A.6 extend to the current setting, largely due to similar properties of the identity and the log links. The main difference is the need to replace the identity link from Section A.6 and the intercept terms absorbing constants from MGFs; a similar phenomenon has been discussed in Section A.3. Due to algebraic similarities, we omit the details and directly state the results.
A.7.1 Logit $Y$ Link and Log $W$ Link Combination with No Interaction Term
@@
$U|A,Z,Y=0\sim\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|Y=0;$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.
@@
Following algebra in Section A.3 and A.6.1, we obtain the following two regression equations which imply Result 12:
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=0]=t(A,Z)$, i.e.
@@
A.7.2 Logit $Y$ Link and Log $W$ Link Combination with Interaction Terms
@@
$U|A,Z,Y=0\sim\underbrace{E[U|A,Z,Y=0]}_{=t(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|Y=0;$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0$ exists for any real value *.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,Y=0]=t(A,Z)$, i.e.
@@
A.7.3 Log $Y$ Link and Logit $W$ Link Combination with No Interaction Term